Suppose that the total market demand for products compresses the demand of two individuals with identical demand equations Qd1=Qd2=Q3=50-25p what is the market demand equation
1
Expert's answer
2021-07-29T12:40:01-0400

"Q_1=50-25P"...Individual 1

"Q_2=50-25P"...Individual 2

The market Demand Equation will remain the same:

"Q_m=50-25P"